We want to calculate the average concentration of a relatively short acting drug in a patient. A bolus of the drug is given quickly via intravenous (IV) and the initial concentration of the drug in the patient is 20 mg/dL. The drug has a half-life of 2 hours. What is the average concentration of the drug in this patients body over 5 half-lives? Assume first-order kinetics.
Answer #1

Ans. First order kinetics-

ln ([A]t / [A]0) = -kt                                -equation 1

where, [A]t = Final concentration

            [A]0 = Initial concentration

            k = rate constant

            t = time of reaction

Also, Half-life of first order reaction, t½ = 0.693 / k

            Or, k = 0.693 / t½                                        - equation 2

Combining equation 1 and 2 –

            ln ([A]t / [A]0) = -(0.693 / t½) t               

            or, ln [A]t – ln [A]0 = -(0.693 / t½) t                    - equation 3

Given-

            Initial [drug], [A]0 = 20 mg/ dL

            Time, t = 5 x (half-life) = 5 x 2 hrs = 10 hrs

Now,

Putting the values in equation 3-

            ln [A]t – ln (20) = -(0.693 / 2 hr ) x 10 hr

            Or, ln [A]t – 2.9957 = - 3.465

            Or, 2.303 log [A]t = - 3.465 + 2.9957 = -0.4693

            Or, log [A]t = -0.4693 / 2.303 = -0.203778

            Or, [A]t = antilog (- 0.203778) = 0.625

Therefore, final [Drug] in patient’s body, [A]t = 0.625 mg/ dL